light travels at a rate of 1.86 times 10 to the fifth power. if the distance from a star to a planet in its orbit is 80 million miles,find how long it takes the light of the star to reach the planet.
Answer by stanbon(75887)

light travels at a rate of 1.86 times 10 to the fifth power. if the distance from a star to a planet in its orbit is 80 million miles,find how long it takes the light of the star to reach the planet.
--------------
time = distance/rate
-----
time = (80*10^6 miles)/(1.86*10^5 miles/sec) = 43*10
= 430.1075 seconds
= 7.167 minutes
-----------------
Cheers,
Stan H.